# Grammatical Error Correction for Low-Resource Languages: The Case of Zarma

### v2

@@ 6.2 Manual Evaluation Results
Given the limitations of AE metrics in low-resource settings, we conducted the ME involving five native Zarma speakers. The evaluators assessed how well each method corrected LG errors, using a scale from 1 to 5, where 1 indicates a poor correction and 5 indicates an excellent correction. The evaluation was performed on a set of 300 sentences from the Feriji Dataset, which were manually corrupted by our annotators to include LG errors. Detailed information about the evaluation protocol and the recruitment of evaluators can be found in Section C. The results are presented in Table 4.
@@ 6.3 Zero-Shot scores
To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset. For that, We gathered a separate set of Zarma texts from books that had not been used during training. We then applied the same corruption script described in Section 4 and asked the annotators to include some LG errors. containing sentences with words and grammatical structures not present in the training data. This evaluation aimed to understand how well the models can handle new or unseen data, which is a common scenario in low-resource language settings. The results of the zero-shot evaluation are presented in Table 5.
@@
GLEU Detection Correction
Rule-based - - 94.17% 81.04% 0 0
Gemma 2b 0.3926 0.481 53.93% 26.52% 0.6 0
MT5-small 0.5018 0.657 73.64% 58.31% 0.8 1
M2M100 0.7683 0.904 92.27% 78.34% 2.7 1.3
@@
As expected, the rule-based method showed a decrease in performance compared to the controlled evaluation. Its detection rate dropped to 94.17% and suggestion accuracy to 81.04%. This decline is due to the method’s reliance on a predefined dictionary and rules, which may not cover new words or grammatical constructions present in the zero-shot dataset. Additionally, the ME scored it with 0 points for both logical errors correction and sentence improvement.
The M2M100 model again demonstrated higher performance among the tested models in the zero-shot setting. It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability. Its detection rate was 92.27% with a suggestion accuracy of 78.34%, showing that it can effectively identify and correct errors in unseen data. Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
The LLM-based models, Gemma 2b and MT5-small, showed reduced performance in the zero-shot evaluation. Gemma 2b’s detection rate was 53.93% with a suggestion accuracy of 26.52%, while MT5-small achieved a detection rate of 73.64% and a suggestion accuracy of 58.31%. In terms of ME, Gemma 2b scored 0.6 for logical errors correction and 0 for sentence improvement, while MT5-small achieved 0.8 and 1, respectively.
Our results suggest that MT models trained on multilingual datasets hold significant potential for improving GEC in low-resource languages. The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks. To validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family. The results of this experiment are detailed in Section A.
7 Potential Applications
Content Creation and Educational Resources
Cultural Preservation and Documentation
8 Conclusion and Future Work
We presented a comparison of rule-based, LLM-based, and MT-based methods for Zarma GEC. The MT-based approach performed best in both controlled and zero-shot scenarios. We confirmed these findings by testing with Bambara language. This research establishes a foundation for future GEC development in other low-resource languages.
9 Limitations
9.1 Rule-Based Method
9.2 LLMs Method
9.3 MT Method
Appendix A Further Experiment with other Languages
After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models. We selected the Bambara language for this experiment because it belongs to a different linguistic family, which allows us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family. We used the Bayelemabaga dataset (Vydrin et al., 2022) for this experiment. The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance. The results are presented in Table 6.
@@ Appendix A Further Experiment with other Languages
The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics. The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%. In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy. The results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly. However, they also emphasize the necessity for continued expanding and diversifying training datasets.
Appendix B Errors Being Addressed
@@
In this section, we explain the types of errors GEC methods address. We categorize the errors into two main types: word-level correction (spellchecking) and context-level correction. The context-level correction is further divided into grammar errors, logical errors, and sentence improvement. Below, we define each error type and provide examples to illustrate them.
B.1 Word-Level Correction (Spellchecking)
@@
Word-Level correction involves identifying and correcting typographical errors in individual words. These errors are usually due to misspellings, incorrect usage of characters, or typographical mistakes.
@@
    • Incorrect: Sintina gaa Irikog na beena da ganda taka.
    • Correct: Sintina gaa Irikoy na beena da ganda taka.
B.2 Context-Level Correction
Context-level correction involves errors that go beyond individual words and affect the overall structure and meaning of the sentence. We categorize these errors into logical errors and sentence improvement.
B.2.1 Logical Errors
Logical errors are inconsistencies in sentence meaning that arise from incorrect verb conjugations, subject-verb agreement, or the misuse of grammatical markers. Unlike grammatical errors, logical errors affect the semantic coherence of a sentence.
@@ B.2.1 Logical Errors
    • Incorrect: Souba, Ay koy Niamey. (The time indicator "Souba" means "tomorrow," but the verb "koy" indicates present tense.)
    • Correct: Souba, Ay ga koy Niamey. (The future tense marker "ga" matches the time indicator "Souba.")
B.2.2 Sentence Improvement
Sentence improvement focuses on enhancing readability, precision, and style. While the sentences may already be grammatically correct, improvements ensure better flow and clarity.
@@ B.2.2 Sentence Improvement
    • Original: I girbi honkuna i tun be.
    • Improved: I ga girbi suba.
Appendix C Evaluators Recruitment Process
To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators. The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language. The evaluators were selected through community outreach in Niamey, Niger. All selected evaluators had completed secondary education or higher.
@@ Appendix C Evaluators Recruitment Process
All evaluators provided informed consent to participate in the study and were compensated for their time and contributions.
Appendix D Training Details
Models Parameters Training Details
QLoRA GPU Used Lr
Gemma 2b 2 billion Applied P100 $2\times 10^{-4}$
MT5-small 300 million Not Applied T4x2 $2\times 10^{-5}$
M2M100 418 million Not Applied P100 $2\times 10^{-5}$
Table 7: Training Settings for the models
Appendix E Other Figures
Figure 2: Images of the different GEC tool interfaces. The rule-based on the left and the other approaches on the right

### v3

@@ 6.3 Manual Evaluation Results
Given the limitations of AE metrics in low-resource settings, we conducted the ME involving five native Zarma speakers. The evaluators assessed how well each method corrected LG errors, using a scale from 1 to 5, where 1 indicates a poor correction and 5 indicates an excellent correction. The evaluation was performed on a set of 300 sentences from the Feriji Dataset, which were manually corrupted by our annotators to include LG errors. Detailed information about the evaluation protocol and the recruitment of evaluators can be found in Section B. The results are presented in Table 5.
@@ 6.4 Zero-Shot Scores
GLEU Detection Correction False Positive
Rule-based 0.2800 0.365 94.17% 81.04% 3.8% 0 0
Gemma 2b 0.3926 0.481 53.93% 26.52% 18.2% 0.6 0
MT5-small 0.5018 0.657 73.64% 58.31% 10.5% 0.8 1
M2M100 0.7683 0.904 92.27% 78.34% 5.1% 2.7 1.3
@@
To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset. For that, we gathered a separate set—150 sentences—of Zarma texts from books that had not been used during training. We then applied the same corruption script and asked the annotators to include some LG errors, containing sentences with words and grammatical structures not present in the training data. This evaluation aimed to understand how well the models can handle new or unseen data, which is a common scenario in low-resource language settings. The results of the zero-shot evaluation are presented in Table 6.
As expected, the rule-based method showed a decrease in performance compared to the controlled evaluation. Its GLEU score dropped to 0.2800 and M² to 0.365, with a detection rate of 94.17%, suggestion accuracy of 81.04%, and a false positive rate of 3.8%. This decline is due to the method's reliance on a predefined dictionary and rules, which may not cover new words or grammatical constructions present in the zero-shot dataset, though it maintains a low rate of incorrect flagging. Additionally, the ME scored it with 0 points for both logical errors correction and sentence improvement.
The M2M100 model again demonstrated higher performance among the tested models in the zero-shot setting. It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability. Its detection rate was 92.27% with a suggestion accuracy of 78.34% and a false positive rate of 5.1%, showing that it can effectively identify and correct errors in unseen data while keeping over-correction low. Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
The LLM-based models, showed reduced performance in the zero-shot evaluation. Gemma 2b's detection rate was 53.93% with a suggestion accuracy of 26.52% and a false positive rate of 18.2%, while MT5-small achieved a detection rate of 73.64%, a suggestion accuracy of 58.31%, and a false positive rate of 10.5%. In terms of ME, Gemma 2b scored 0.6 for logical errors correction and 0 for sentence improvement, while MT5-small achieved 0.8 and 1, respectively. These higher false positive rates indicate a greater tendency to misidentify correct words as errors in unseen data.
Our results suggest that MT models trained on multilingual datasets hold significant potential for improving GEC in low-resource languages. The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks. The lower accuracy of LLMs likely arises from their lack of initial pre-training on Zarma and the use of quantization during training, which may impact performance due to limited computational resources in our low-resource setting. In contrast, M2M100 performed well despite these constraints. We hypothesize that in higher computational resource settings, LLMs could outperform all other methods, leveraging their capacity for nuanced language understanding when adequately pre-trained and resourced.
Nonetheless, to validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family. The results of this experiment are detailed in Section 7.
7 Further Experiment with other Languages
After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models. We selected the Bambara language for this experiment because it belongs to a different linguistic family, which allows us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family. We used the Bayelemabaga dataset (Vydrin et al., 2022)—split into 80% 10% 10%—for this experiment. The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance. The results are presented in Table 7.
@@ 7 Further Experiment with other Languages
The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics. The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%. In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy. The results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly. However, they also emphasize the necessity for continued expanding and diversifying training datasets and pre-training for LLMs.
8 Conclusion and Future Work
We presented a comparison of rule-based, LLM-based, and MT-based methods for Zarma GEC. Our results provide answers to our research questions. Regarding RQ1, we found that MT models (M2M100) outperform both rule-based systems and LLMs like Gemma and MT5 in Zarma GEC. For RQ2, we observed that rule-based methods excel at simple spelling correction but fail at context-level errors, whereas MT models handle both effectively. Finally, for RQ3, our experiments with Bambara confirm that the MT-based approach is replicable and effective for other low-resource West African languages. This research establishes a foundation for future GEC development in other low-resource languages. The approaches covered in this study are starting point references and can be further improved using sophisticated methods.
Models Parameters Training Details
QLoRA Epochs Batch Size GPU Used Lr
Gemma 2b 2 billion Applied 2 8 P100 $2\times 10^{-4}$
MT5-small 300 million Not Applied 10 16 T4x2 $2\times 10^{-5}$
M2M100 418 million Not Applied 2 32 P100 $2\times 10^{-5}$
Table 8: Training settings for the models, including epochs, batch size, and optimizer details.
Limitations
Rule-Based Method
LLMs Method
MT Method
Synthetic Data Generation
Our synthetic data generation process—while designed to mimic common human errors based on an initial study with native speakers (as detailed in Section 4)—has some limitations. The corruption script may not capture the full nuanced distribution of errors that humans make—particularly complex grammatical or semantic errors that are not easily programmable by rule-based mechanisms. This could lead to models being better prepared for certain types of synthetic errors than for the true wide range of errors encountered in real-world Zarma texts. Consequently, performance on naturally occurring errors might differ from that observed on our synthetic test sets.
Acknowledgments
This research was partially supported by the computer science department of Ashesi University, and the Cite de l'Innovation of Niger.
Appendix A Errors Being Addressed
@@
In this section, we explain the types of errors GEC methods address. We categorize the errors into two main types: word-level correction (spellchecking) and context-level correction. The context-level correction is further divided into logical errors and sentence improvement. Below, we define each error type and provide examples in Zarma with English translations to illustrate them.
A.1 Word-Level Correction
@@
Word-Level correction involves identifying and correcting typographical errors in individual words. These errors are usually due to misspellings, incorrect usage of characters, or typographical mistakes. Such errors typically do not alter the grammatical structure of the sentence but affect readability and lexical accuracy.
@@
    • Incorrect Zarma: Sintina gaa Irikog na beena da ganda taka.
    • English Translation (Incorrect): In the beginning God [misspelled form of God] created the heaven and the earth.
    • Correct Zarma: Sintina gaa Irikoy na beena da ganda taka.
    • English Translation (Correct): In the beginning God created the heaven and the earth.
    • Explanation: The word ``Irikog'' is a misspelling of ``Irikoy'' (God). Spellchecking corrects this specific lexical error.
A.2 Context-Level Correction
Context-level correction involves errors that go beyond individual words and affect the overall structure, coherence, and meaning of the sentence. These errors require an understanding of the surrounding context to identify and rectify. We categorize these errors into logical errors and sentence improvement.
A.2.1 Logical Errors
Logical errors are inconsistencies in sentence meaning often caused by incorrect verb conjugations, subject-verb agreement issues, or the misuse of grammatical markers such as tense or aspect markers. These errors affect the semantic coherence of the sentence and require contextual understanding to correct.
@@ A.2.1 Logical Errors
    • Incorrect Zarma: Souba, Ay koy Niamey.
    • English Translation (Incorrect): Tomorrow, I go to Niamey.
    • Correct Zarma: Souba, Ay ga koy Niamey.
    • English Translation (Correct): Tomorrow, I will go Niamey.
    • Explanation: The time indicator ``Souba'' means ``tomorrow,'' which requires a future tense. The original sentence incorrectly uses ``koy'' (go/am going), which is present tense. The correction introduces the future tense marker ``ga'' before the verb ``koy'' to align the verb tense with the future time reference, resolving the logical inconsistency.
A.2.2 Sentence Improvement
Sentence improvement focuses on enhancing the overall readability, precision, naturalness, or stylistic quality of a sentence. While the original sentence may already be grammatically correct, improvements aim to make it flow better, be more concise, or adhere more closely to common usage patterns.
@@ A.2.2 Sentence Improvement
    • Original Zarma: Albarka ma te ni se teŋ kaŋ ni goy.
    • English Translation: Blessings be to you for the work that you did.
    • Improved Zarma: Ni goyom albarka.
    • English Translation: Blessings on your work / Your work is blessed.
    • Explanation: The original sentence is grammatically correct but might be perceived as slightly formal or verbose in some contexts. The improved version offers a more concise and idiomatic expression common in Zarma for acknowledging good work.
Appendix B Evaluators Recruitment Process
To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators, who also assisted in developing the corruption script for synthetic data (Section 4). The primary criterion for selection was their fluency in Zarma—their ability to read and write proficiently in the language. The evaluators were selected through community outreach in Niamey, Niger. All selected evaluators had completed secondary education or higher.
@@ Appendix B Evaluators Recruitment Process
All evaluators gave informed consent to participate in the study and were compensated for their time and contributions.
Appendix C Potential Applications
Content Creation and Educational Resources
Cultural Preservation and Documentation